OpenStudy (anonymous):

Given equations A and B as 2/5x+y=12 and 5/2y-x=6 , respectively, which expression will eliminate the variable x?

OpenStudy (anonymous):

@dan815

OpenStudy (dan815):

what are the expressions

OpenStudy (anonymous):

5/2 A +2/5 B 5/2 A − B A + 2/5 B 5/2A − 2/5 B 25/10 A + 6/10 B

OpenStudy (anonymous):

and yes dan

OpenStudy (anonymous):

hurrrrrrry pplease

OpenStudy (dan815):

5/2 A − B will leave you with a 5/2 Y in the first equation so it will cancel with equation B

OpenStudy (anonymous):

thank you !!!!!!!!!!!! <3

OpenStudy (anonymous):

Which values of x and y will satisfy y = 5x and 12x + 9y = 60? x = 1, y = 19 x = 35, y = 20/19 x = 1/20 y = 1/90 x = 20/19 y = 100/19 x = 19, y = 50/19

OpenStudy (dan815):

sub the 5x in to u for equation 2

OpenStudy (dan815):

12x+9*5x=60 57x=60 x=60/57=20/19

OpenStudy (anonymous):

-_-

OpenStudy (dan815):

2nd last choice stop using me now =[ u hut meh feels

OpenStudy (anonymous):

im not using u im asking you :C

OpenStudy (dan815):

lol :)

OpenStudy (anonymous):

The difference of two numbers, a and b, is 21. The difference of 5 times a and two times b is 18. What are the values of a and b? 4, 1 -19, 40 20, 1 30, 9 -8, -29 :* your like a huge help tho i need to pass this test and your so nice to help

OpenStudy (dan815):

a-b=21 5a-2b=18

OpenStudy (anonymous):

b?

OpenStudy (dan815):

no let me see

OpenStudy (anonymous):

yay thats what i got <3

OpenStudy (dan815):

:)